OpenStudy (anonymous):

Solve these proportions. x/3 = x + 2/2 x=

OpenStudy (anonymous):

\[x/3=(x+2)/2\] Multiply both sides by 2 \[2x/3=x+2\] Multiply both sides by 3. \[2x=(x+2)*3\] So now you have 2x=3x+6 Subtract 6 from both sides So now you have 2x-6=3x Subtract 2x from both sides So now you have -6=3x-2x This becomes -6=x

OpenStudy (anonymous):

they want it like this = ? sqrt ?

OpenStudy (anonymous):

That doesn't make much sense to me. But you could write -6 as \[-1\sqrt{36}\]